OpenStudy (anonymous):

Let f(x) = x + 8 and g(x) = x2 – 6x – 7. Find f(g(2)).

OpenStudy (anonymous):

g(2)=(2)^2-6(2)-7 g(2)=4-12-7 g(2)=4-19 g(2)=-15

OpenStudy (anonymous):

f(g(2))=f(-15) f(g(2))=-15+8 f(g(2))=-7

OpenStudy (anonymous):

got it